Here is some data on the economy of a certain country. Use this data for all three parts of question 1.

                             

                            M1 Money Supply:   $190 Billion

                                             Real GDP:   $765 Billion

   Velocity of M1 Money Supply:   4.3

Question 1a: Right now, what is the rate of inflation in that country? How can you tell?

Question 1b: If they want to change the inflation rate to about 2% (which the Fed says is ideal), can they do so by changing the Money Supply? Assuming the velocity stays the same, about how much should the Money Supply be to get to the inflation rate to 2%

Question 1c: Give one example of action the Central Bank of that country could take to reduce the money supply. Explain how your example would work. For this question, assume that the Central Bank of that country has tools for Monetary Policy similar to the tools used by the F.O.M.C. in the USA. You don’t need specific numbers to answer this, just a clear explanation.
Answer #1

Consider the quantity theory of money supply as given in the following form -

PY

Given -

M = 190 billion

V = 4.3

Y = 765 billion

We get inflation/price level as. -

P=\frac{MV}{Y}

\Rightarrow P=\frac{(190)(4.3)}{765}

\Rightarrow P=1.067

If P = 2, V = 4.3 and Y = 765, we get -

M=\frac{PY}{V}

\Rightarrow M=\frac{(2)(765)}{4.3}

\Rightarrow M=356

One way for the Central Bank to increase the money supply in the economy is to reduce the discount rate. This is the rate at which the central bank lends money to the commercial banks. If this rate is reduced, then the commercial banks can borrow money from the central bank at a lower interest rate, which will reduce the cost of borrowing for the commercial banks, in which case they will borrow more and advance more loans to the public, which will increase the money supply in the economy.